OpenStudy (anonymous):

Solve the following problems using the sequence: 5, 13, 21, 29, 37, ...... a. Use words to write a rule for the sequence. b. Use the rule you wrote in the previous part to find the next three terms of the sequence. Be sure to show all your steps.

OpenStudy (mathmath333):

try to find the difference between the number and the next number

OpenStudy (anonymous):

a) find the next term by adding eight to the previous term. 5 + 8 = 13 + 8 = 21 + 8 = 29 + 8 = 37 b) 45, 53, 61 5 + 8 = 13 + 8 = 21 + 8 = 29 + 8 = 37 + 8 = 45 + 8 = 53 + 8 = 61 is this right ?